from the collection of lotus flowers, 1/3rd offered to shiva, 1/5th to vishnu, 1/6th to the sun and 1/4th to the saraswati.then,there were 6 remaining.Find the toal number of lotus flowers.
Answer by stanbon(75887)

from the collection of lotus flowers, 1/3rd offered to shiva, 1/5th to vishnu, 1/6th to the sun and 1/4th to the saraswati.then,there were 6 remaining.Find the toal number of lotus flowers.
-------------------------
Let "x" be the total number of lotus flowers:
Equation:
x - x/3 -x/5 -x/6 - x/4 = 6
------
Multiply thru by 60 to get:
60x - 20x - 12x -10x - 15x = 360
-----
60x - 57x = 360
3x = 360
x = 120 (# of lotus flowers)
================================
Cheers,
Stan H.
=============